The Chinese Alligator (Alligator sinensis; above) is a critically endangered species living in a restricted reserve area in the Yangtze River in China’s Anhui province. (Captive populations are doing much better; where individuals may live up to 70 years).

(3 pts) After spraying an insecticide the number of individuals of each genotype that survived is given below. Calculate the absolute fitness, and relative fitness of each genotype. Round up the answers to three decimal points.

Genotype

Number of individuals survived

AA

50

Aa

18

aa

15
Answer #1

From the given data,

Assuming that the population size is 500, the number of alleles is = 1000

The frequency of A = frequency of AA + 1/2 frequency of Aa = 50/83 + 1/2 (18/83) = 0.71

The frequency of a =  frequency of aa + 1/2 frequency of Aa = 15/83 + 1/2 (18/83) = 0.288

The expected frequencies are,

AA = 0.71*0.71 = 0.5041; number of alleles is = 0.504*1000 = 504

Aa = 0.71* 0.288 = 0.204; number of alleles is = 0.204*1000 = 204

aa = 0.288*0.288 = 0.082; number of alleles is = 0.288*1000 = 288

The absolute fitness is,

AA: 50/504 = 0.099

Aa: 18/204 = 0.088

aa: 15/288 = 0.052

The relative fitness is,

AA: 0.099/0.099 = 1

Aa: 0.088/0.099 = 0.88

aa: 0.052/0.099 = 0.525